# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Uncertain Dynamics

### v2

@@ VI Simulation and Experimental Results
VI-A Planning in an Obstacle-Dense Environment
To evaluate the effectiveness of the VF-LPC approach, we compare it with advanced planning approaches in a scenario containing static obstacles; see Fig. 7. We compare VF-LPC with the commonly used planning approaches, such as Hybrid A*[38], RRT*-CFS[39], Timed Elastic Bands (TEB)[40] and OBTPAP[41] which iteratively solves the time-optimal motion planning problem. In these tests, we use the analytic kinematic vehicle model with the state variable $[X,Y,\psi]$, representing the $X$-, $Y$-coordinates and the heading, respectively. All the approaches are performed in MATLAB 2023a.
@@ VI-B1 Simulation settings
All the comparison methods use the system dynamics model (43) and solve the nonlinear optimization problem with the IPOPT solver[45] under CasADi[46]. The vehicle dynamics are described using a “bicycle” model [47] as follows:
@@
For VF-LPC, we employed the approach outlined in [21] to conduct system identification by collecting vehicle motion data from the CarSim solver. Finally, a linear time-invariant system model (4) is generated for the VF-LPC algorithm, where $n_{\mathcal{K}}=10$ and $n_{u}=2$. To construct an error model for facilitating subsequent algorithm design, we subtract the desired state from the current state, i.e, $\tilde{x}=x-x_{r}$ and the desired control from system control, i.e., $\tilde{u}=u-u_{r}=[a_{x},\delta_{f}]^{\top}$.
@@ VI-C Model Uncertainties Learning
[…]
The algorithm was simulated in a racing track road and the desired speed $v_{d}$ is set to be $6\ \rm{m/s}$. Note that we collect motion data of the real vehicle (right half of Fig. 4) to train for an offline nominal deep Koopman model with $n_{\rho_{e}}=6$ and $x=[X,Y,\psi,v_{x},v_{y},\omega]^{\top}\in\mathbb{R}^{6}$. The exact vehicle parameters are: $m=1257\rm{kg}$, $I_{z}=1524.9\rm{kg}\cdot\rm{m}^{2}$, $C_{af}=8790\rm{N/rad}$, and $C_{ar}=30400\rm{N/rad}$. The iid process noise $w_{k}\in\mathbb{R}^{6}$ with zero mean and variance $\sigma=0.002$ is added to the vehicle dynamics. Specifically, $g(x,u)$ maps $[x,u]^{\top}$ to $[v_{x},v_{y},\omega]^{\top}$.
@@ VI-D Validation on Unmanned Aerial Vehicles
Firstly, we conduct data-driven modeling of the quadrotor UAV using deep Koopman operators[21], followed by employing the VF-LPC to generate the robot’s maneuvers in a 3D environment, as illustrated in Fig. 10. The state variables of the training data consist of the XYZ coordinates and velocities ($v_{x},v_{y},v_{z}$) of the quadrotor UAV, with control variables of acceleration in the XYZ directions. A total of $148$ data, each comprising approximately $4500$ samples, were collected randomly in the state and control space for training and model validation purposes. The $6$-dimensional state vector $[X,Y,Z,v_{x},v_{y},v_{z}]^{\top}$ was processed through an encode layer, resulting in a final system model of $18$ state variables and $3$ control variables (i.e., $n_{\mathcal{K}}=18,n_{u}=3$). Then the obtained model is incorporated into VF-LPC, which focuses on planning 3-axis accelerations for attitude control. Not only the VF-LPC approach can avoid multiple static obstacles. Under suddenly appearing obstacles (see Fig. 10), the VF-LPC approach is also capable of handling them, and the average speed of the vehicle reaches $1\rm{m/s}$.

### v4

@@ VI Simulation and Experimental Results
VI-A Path Planning in An Obstacle-Dense Environment
To evaluate the effectiveness of the VF-LPC approach, we compare it with advanced planning approaches in a scenario containing static obstacles; see Fig. 7. We compare VF-LPC with the commonly used planning approaches, such as Hybrid A*[40], RRT*-CFS[41], Timed Elastic Bands (TEB)[42] and OBTPAP[43] which iteratively solves the time-optimal motion planning problem. In these tests, we use the analytic kinematic vehicle model with the state variable $X,Y,\psi]$, representing the $X$-, $Y$-coordinates and the heading, respectively. All the approaches are performed in MATLAB 2023a.
@@ VI-B1 Simulation settings
All the comparison methods utilize the system dynamics model (49) and solve the nonlinear optimization problem using the IPOPT solver[47] within the CasADi framework[48], which describes the vehicle dynamics as follows:
@@
For VF-LPC, we employed the approach outlined in [21] to conduct system identification by collecting vehicle motion data from the CarSim solver. Finally, a linear time-invariant system model (4) is generated for the VF-LPC algorithm, where $n_{\mathcal{K}}=10$ and $n_{u}=2$. To construct an error model for facilitating subsequent algorithm design, we subtract the desired state from the current state, i.e, $\tilde{x}=x-x_{r}$ and the desired control from system control, i.e., $\tilde{u}=[a_{x},\delta_{f}]^{\top}$.
@@ VI-C Model Uncertainties Learning
The uncertainty in the dynamics of a system negatively affects the planning and control performance of robots. To eliminate the influence, a quantified sparse GP [9] is utilized to compensate for the difference between the offline-trained deep Koopman model and the exact vehicle model. The simulated dynamical model (49), reveals that the vehicle states $v_{x},v_{y},\omega$ and $\rho_{e}(x_{k})$ are primarily affected by dynamic parameters. Therefore, we further derive $B_{d}=[\mathbf{0}_{3}\ I_{3}\ \mathbf{0}_{3\times n_{\rho_{e}}}]^{\top}$, and $g(x,u)=g(v_{x},v_{y},\omega,\rho_{e}^{\top},a_{x},\delta_{f}):\mathbb{R}^{n_{\rho_{e}}+5}\rightarrow\mathbb{R}^{3}$.
The algorithm was simulated in a racing track road and the desired speed $v_{d}$ is set to be $6\ \rm{m/s}$. Note that we collect motion data of the real vehicle (right half of Fig. 4) to train for an offline nominal deep Koopman model with $n_{\rho_{e}}=6$ and $x=[X,Y,\psi,v_{x},v_{y},\omega]^{\top}\in\mathbb{R}^{6}$. The exact vehicle parameters are: $m=1257\rm{kg}$, $I_{z}=1524.9\rm{kg}\cdot\rm{m}^{2}$, $C_{af}=8790\rm{N/rad}$, and $C_{ar}=30400\rm{N/rad}$. The iid process noise $w_{k}\in\mathbb{R}^{6}$ with zero mean and variance $\sigma=0.002$ is added to the vehicle dynamics. Specifically, $g(x,u)$ maps $x,u]^{\top}$ to $v_{x},v_{y},\omega]^{\top}$.
@@ VI-D Validation on Unmanned Aerial Vehicles
Firstly, we conduct data-driven modeling of the quadrotor UAV using deep Koopman operators[21], followed by employing the VF-LPC to generate the robot’s maneuvers in a 3D environment, as illustrated in Fig. 10. The state variables of the training data consist of the XYZ coordinates and velocities ($v_{x},v_{y},v_{z}$) of the quadrotor UAV, with control variables of acceleration in the XYZ directions. A total of $148$ data, each comprising approximately $4500$ samples, were collected randomly in the state and control space for training and model validation purposes. The $6$-dimensional state vector $X,Y,Z,v_{x},v_{y},v_{z}]^{\top}$ was processed through an encode layer, resulting in a final system model of $18$ state variables and $3$ control variables (i.e., $n_{\mathcal{K}}=18,n_{u}=3$). Then the obtained model is incorporated into VF-LPC, which focuses on planning 3-axis accelerations for attitude control. Not only the VF-LPC approach can avoid multiple static obstacles. Under suddenly appearing obstacles (see Fig. 10), the VF-LPC approach is also capable of handling them, and the average speed of the vehicle reaches $1\rm{m/s}$.